A bond with an annual coupon rate of 5.1% sells for $985. What is the bond’s current yield? (Round your answer to 2 decimal places.)
Answer #1

Annual coupon=1000*5.1%=$51

Hence current yield=Annual coupon/Current price

=51/985

which is equal to

=5.18%(Approx).